OpenStudy (anonymous):

Jim's uncle is 4 times as old as his nephew Jim. In 10 years, Jim's uncle age will be 20 years more than twice Jim's age. How old is Jim right now?

TheSmartOne (thesmartone):

Jim's Uncle is U Jim is J

OpenStudy (anonymous):

nah Jim is x and the uncle is 4x

TheSmartOne (thesmartone):

So U=4J

TheSmartOne (thesmartone):

and (U+10)=20+2J

TheSmartOne (thesmartone):

So plug in U=4J into U+10=20+2J

TheSmartOne (thesmartone):

4J+10=20+2J 2J=10 J=5

TheSmartOne (thesmartone):

So Jim is 5 years old.

OpenStudy (anonymous):

Solve the following for u and n, uncle and nephew respectively.\[\{u=4 n,u+10=2 (n+10)+20\} \]\[\{n=15,u=60\} \]

TheSmartOne (thesmartone):

Oh yea I made a mistake... Good Job @robtobey

OpenStudy (anonymous):

Thank you for the medal.